OpenStudy (anonymous):

3/x-2 = 2/x-3 -13 0 5 13

OpenStudy (anonymous):

cross-multiply 3(x-3) = 2(x-2) 3x - 9 = 2x - 4 can u do the rest?

OpenStudy (anonymous):

would it be 5?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

cool. thanks!!